\begin{thebibliography}{99}


%\cite{Brezin:1977sv}
\bibitem{Brezin:1977sv} 
  E.~Brezin, C.~Itzykson, G.~Parisi and J.~B.~Zuber,
  %``Planar Diagrams,''
  Commun.\ Math.\ Phys.\  {\bf 59}, 35 (1978).
%  doi:10.1007/BF01614153
  %%CITATION = doi:10.1007/BF01614153;%%
  %1165 citations counted in INSPIRE as of 08 Aug 2016


%\cite{Shimamune:1981qf}
\bibitem{Shimamune:1981qf} 
  Y.~Shimamune,
%  ``On The Phase Structure Of Large N Matrix Models And Gauge Models,''
  Phys.\ Lett.\ B {\bf 108}, 407 (1982).
  %%CITATION = PHLTA,B108,407;%%




%\cite{O'Connor:2007ea}
\bibitem{O'Connor:2007ea} 
  D.~O'Connor and C.~Saemann,
%  ``Fuzzy Scalar Field Theory as a Multitrace Matrix Model,''
  JHEP {\bf 0708}, 066 (2007).
%  [arXiv:0706.2493 [hep-th]].
  %%CITATION = ARXIV:0706.2493;%%

%\cite{Saemann:2014pca}
\bibitem{Saemann:2014pca} 
  C.~Saemann,
  %``Bootstrapping Fuzzy Scalar Field Theory,''
  JHEP {\bf 1504}, 044 (2015).
%  doi:10.1007/JHEP04(2015)044
%  [arXiv:1412.6255 [hep-th]].
  %%CITATION = doi:10.1007/JHEP04(2015)044;%%
  %5 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Rea:2015wta}
\bibitem{Rea:2015wta} 
  S.~Rea and C.~Saemann,
  %``The Phase Diagram of Scalar Field Theory on the Fuzzy Disc,''
  JHEP {\bf 1511}, 115 (2015).
%  doi:10.1007/JHEP11(2015)115
%  [arXiv:1507.05978 [hep-th]].
  %%CITATION = doi:10.1007/JHEP11(2015)115;%%
  %2 citations counted in INSPIRE as of 08 Aug 2016




\bibitem{Hoppe} 
J. Hoppe, MIT Ph.D. Thesis, (1982). 


%\cite{Madore:1991bw}
\bibitem{Madore:1991bw} 
  J.~Madore,
  %``The Fuzzy sphere,''
  Class.\ Quant.\ Grav.\  {\bf 9}, 69 (1992).
%  doi:10.1088/0264-9381/9/1/008
  %%CITATION = doi:10.1088/0264-9381/9/1/008;%%
  %496 citations counted in INSPIRE as of 08 Aug 2016


%\cite{Lizzi:2003ru}
\bibitem{Lizzi:2003ru} 
  F.~Lizzi, P.~Vitale and A.~Zampini,
  %``The Fuzzy disc,''
  JHEP {\bf 0308}, 057 (2003).
%  doi:10.1088/1126-6708/2003/08/057
%  [hep-th/0306247].
  %%CITATION = doi:10.1088/1126-6708/2003/08/057;%%
  %49 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Onsager:1943jn}
\bibitem{Onsager:1943jn} 
  L.~Onsager,
  %``Crystal statistics. 1. A Two-dimensional model with an order disorder transition,''
  Phys.\ Rev.\  {\bf 65}, 117 (1944).
 % doi:10.1103/PhysRev.65.117
  %%CITATION = doi:10.1103/PhysRev.65.117;%%
  %589 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Myers:1999ps}
\bibitem{Myers:1999ps} 
  R.~C.~Myers,
  %``Dielectric branes,''
  JHEP {\bf 9912}, 022 (1999).
%  doi:10.1088/1126-6708/1999/12/022
%  [hep-th/9910053].
  %%CITATION = doi:10.1088/1126-6708/1999/12/022;%%
  %1139 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Azuma:2004zq}
\bibitem{Azuma:2004zq} 
  T.~Azuma, S.~Bal, K.~Nagao and J.~Nishimura,
  %``Nonperturbative studies of fuzzy spheres in a matrix model with the Chern-Simons term,''
  JHEP {\bf 0405}, 005 (2004).
%  doi:10.1088/1126-6708/2004/05/005
%  [hep-th/0401038].
  %%CITATION = doi:10.1088/1126-6708/2004/05/005;%%
  %84 citations counted in INSPIRE as of 08 Aug 2016



%\cite{Alekseev:2000fd}
\bibitem{Alekseev:2000fd} 
  A.~Y.~Alekseev, A.~Recknagel and V.~Schomerus,
  %``Brane dynamics in background fluxes and noncommutative geometry,''
  JHEP {\bf 0005}, 010 (2000).
%  doi:10.1088/1126-6708/2000/05/010
%  [hep-th/0003187].
  %%CITATION = doi:10.1088/1126-6708/2000/05/010;%%
  %196 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Iso:2001mg}
\bibitem{Iso:2001mg} 
  S.~Iso, Y.~Kimura, K.~Tanaka and K.~Wakatsuki,
  %``Noncommutative gauge theory on fuzzy sphere from matrix model,''
  Nucl.\ Phys.\ B {\bf 604}, 121 (2001).
%  doi:10.1016/S0550-3213(01)00173-0
%  [hep-th/0101102].
  %%CITATION = doi:10.1016/S0550-3213(01)00173-0;%%
  %178 citations counted in INSPIRE as of 08 Aug 2016

%\cite{CarowWatamura:1998jn}
\bibitem{CarowWatamura:1998jn} 
  U.~Carow-Watamura and S.~Watamura,
  %``Noncommutative geometry and gauge theory on fuzzy sphere,''
  Commun.\ Math.\ Phys.\  {\bf 212}, 395 (2000).
%  doi:10.1007/s002200000213
%  [hep-th/9801195].
  %%CITATION = doi:10.1007/s002200000213;%%
  %141 citations counted in INSPIRE as of 08 Aug 2016




%\cite{Steinacker:2005wj}
\bibitem{Steinacker:2005wj} 
  H.~Steinacker,
  %``A Non-perturbative approach to non-commutative scalar field theory,''
  JHEP {\bf 0503}, 075 (2005).
%  doi:10.1088/1126-6708/2005/03/075
%  [hep-th/0501174].
  %%CITATION = doi:10.1088/1126-6708/2005/03/075;%%
  %42 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Nair:2011ux}
\bibitem{Nair:2011ux} 
  V.~P.~Nair, A.~P.~Polychronakos and J.~Tekel,
  %``Fuzzy spaces and new random matrix ensembles,''
  Phys.\ Rev.\ D {\bf 85}, 045021 (2012).
%  doi:10.1103/PhysRevD.85.045021
%  [arXiv:1109.3349 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.85.045021;%%
  %12 citations counted in INSPIRE as of 08 Aug 2016



%\cite{Polychronakos:2013nca}
\bibitem{Polychronakos:2013nca} 
  A.~P.~Polychronakos,
  %``Effective action and phase transitions of scalar field on the fuzzy sphere,''
  Phys.\ Rev.\ D {\bf 88}, 065010 (2013).
%  doi:10.1103/PhysRevD.88.065010
%  [arXiv:1306.6645 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.88.065010;%%
  %16 citations counted in INSPIRE as of 08 Aug 2016

%\cite{Tekel:2013vz}
\bibitem{Tekel:2013vz} 
  J.~Tekel,
  %``Random matrix approach to scalar fields on fuzzy spaces,''
  Phys.\ Rev.\ D {\bf 87}, no. 8, 085015 (2013).
%  doi:10.1103/PhysRevD.87.085015
%  [arXiv:1301.2154 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.87.085015;%%
  %9 citations counted in INSPIRE as of 08 Aug 2016


%\cite{Steinacker:2003sd}
\bibitem{Steinacker:2003sd} 
  H.~Steinacker,
%  ``Quantized gauge theory on the fuzzy sphere as random matrix model,''
  Nucl.\ Phys.\ B {\bf 679}, 66 (2004).
%  doi:10.1016/j.nuclphysb.2003.12.005
%  [hep-th/0307075].
  %%CITATION = doi:10.1016/j.nuclphysb.2003.12.005;%%
  %118 citations counted in INSPIRE as of 07 Aug 2016


%\cite{Connes:1996gi}
\bibitem{Connes:1996gi} 
  A.~Connes,
  %``Gravity coupled with matter and foundation of noncommutative geometry,''
  Commun.\ Math.\ Phys.\  {\bf 182}, 155 (1996).
%  doi:10.1007/BF02506388
%  [hep-th/9603053].
  %%CITATION = doi:10.1007/BF02506388;%%
  %377 citations counted in INSPIRE as of 09 Aug 2016




%\cite{Ishibashi:1996xs}
\bibitem{Ishibashi:1996xs} 
  N.~Ishibashi, H.~Kawai, Y.~Kitazawa and A.~Tsuchiya,
%  ``A Large N reduced model as superstring,''
  Nucl.\ Phys.\ B {\bf 498}, 467 (1997).
%  doi:10.1016/S0550-3213(97)00290-3
%  [hep-th/9612115].
  %%CITATION = doi:10.1016/S0550-3213(97)00290-3;%%
  %858 citations counted in INSPIRE as of 11 Jun 2016

%\cite{Ydri:2016osu}
\bibitem{Ydri:2016osu} 
  B.~Ydri, R.~Ahlam and R.~Khaled,
  %``Emergent fuzzy geometry and fuzzy physics in $4$ dimensions,''
  arXiv:1607.08296 [hep-th].
  %%CITATION = ARXIV:1607.08296;%%

%\cite{Ydri:2015zsa}
\bibitem{Ydri:2015zsa} 
  B.~Ydri, A.~Rouag and K.~Ramda,
  %``Emergent geometry from random multitrace matrix models,''
  Phys.\ Rev.\ D {\bf 93}, no. 6, 065055 (2016).
%  doi:10.1103/PhysRevD.93.065055
%  [arXiv:1509.03572 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.93.065055;%%
  %3 citations counted in INSPIRE as of 07 Aug 2016

%\cite{Ydri:2015vba}
\bibitem{Ydri:2015vba} 
  B.~Ydri, K.~Ramda and A.~Rouag,
  %``Phase diagrams of the multitrace quartic matrix models of noncommutative $\Phi^4$ theory,''
  Phys.\ Rev.\ D {\bf 93}, no. 6, 065056 (2016).
%  doi:10.1103/PhysRevD.93.065056
%  [arXiv:1509.03726 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.93.065056;%%
  %4 citations counted in INSPIRE as of 08 Aug 2016


%\cite{Ydri_in_preparation}
%\bibitem{Ydri_in_preparation} 
%  B.~Ydri,C.~Soudani, A.~Rouag, in preparation.




\end{thebibliography}

\documentclass[
twocolumn,secnumarabic
,amssymb, amsmath,nobibnotes, aps, prl]{revtex4}
\usepackage[dvips]{graphicx}
\usepackage{tabularx}
\usepackage{bm}%
\expandafter\ifx\csname package@font\endcsname\relax\else
 \expandafter\expandafter
 \expandafter\usepackage
 \expandafter\expandafter
 \expandafter{\csname package@font\endcsname}%
\fi

\begin{document}

\title{The multitrace matrix model: An alternative to Connes NCG and IKKT model}

\author{Badis Ydri}
\email{ydri@stp.dias.ie, bydri@ictp.it}
\affiliation{Department of Physics, Badji-Mokhtar Annaba University,\\
 Annaba, Algeria.}

\begin{abstract} 
We present a new multitrace matrix model, which is a generalization of the real quartic one matrix model, exhibiting dynamical emergence of a fuzzy  two-sphere and its non-commutative gauge theory. This provides a novel and a much simpler alternative to Connes non-commutative geometry and to the IKKT matrix model for emergent geometry in two dimensions.
\end{abstract}

\maketitle
The real quartic multitrace model \cite{Brezin:1977sv}
\begin{eqnarray}
V_0&=&{B}Tr M^2+{C}Tr M^4,
\end{eqnarray}
has two stables phases: $i)$ the disordered (symmetric one-cut) phase and $2)$ the non-uniform ordered (two-cut) phase  with the transition being identified of third order. The uniform ordered (Ising asymmetric one-cut) phase is metastable in this theory \cite{Shimamune:1981qf}. It was discovered in  \cite{Ydri:2015vba,Ydri:2015zsa} that the Ising phase becomes stable if we add to $V_0$,  with a particular choice of the parameters $D$, $B^{'}$, $C^{'}$, $D^{'}$, $A^{'}$..., the quartic multitrace model
\begin{eqnarray}
V_1&=&D(Tr M^2)^2+B^{'} (Tr M)^2+C^{'} Tr M Tr M^3\nonumber\\
&+&D^{'}(Tr M)^4+A^{'}Tr M^2 (Tr M)^2+....\label{fundamental1}
\end{eqnarray}
 These are the most general terms consistent with the multitrace expansion of the kinetic term of non-commutative $\Phi^4$ theory at quartic order.  For example, the coefficients on the fuzzy sphere \cite{Hoppe,Madore:1991bw} were calculated in \cite{O'Connor:2007ea,Saemann:2014pca} and on the fuzzy disc \cite{Lizzi:2003ru} were calculated in \cite{Rea:2015wta}. Another set of parameters which give a stable Ising phase are given by \cite{Ydri:2015zsa}
 
\begin{eqnarray}
D=\frac{3N}{4}~,~B^{'}=\frac{\sqrt{N}}{2}~,~C^{'}=-N~,~ D^{'}=0~,~A^{'}=0.\label{our_case}
\end{eqnarray}
This can be explicitly verified by computing the critical exponents across the Ising transition line and the Wigner semi-circle law near in the perturbative regime \cite{Ydri:2015zsa}. The critical exponents are found to be consistent with the Onsager values \cite{Onsager:1943jn} suggesting that the dimension of the underlying space is two and that the above matrix model in this phase is in the Ising universality class. The Wigner semi-circle law is determined by the free propagator and thus by the metric aspects of the underlying space \cite{Steinacker:2005wj,Nair:2011ux, Polychronakos:2013nca,Tekel:2013vz}.
The operator $ TrM TrM^3$ is the crucial ingredient in stabilizing the Ising phase. Indeed, given hindsight, a simpler set of parameters is given simply by  $C^{'}=-N$ while setting all the other parameters to zero. 

The configuration $M=a{\bf 1}_{2N}$, for our case (\ref{our_case}), is a solution of the equation of motion iff
\begin{eqnarray}
a&=&0 ({\rm disordered}),\nonumber\\
a^2&=&-\frac{\tilde{B}+1}{\sqrt{N}(2\tilde{C}-1)}({\rm uniform~ordered}).
\end{eqnarray}
The large $N$ scaling of the various coupling constants is given by 
\begin{eqnarray}
\tilde{B}=\frac{B}{N^{3/2}}~,~\tilde{C}=\frac{C}{N^2}.
\end{eqnarray}
Stability requires that $\tilde{B}\leq -1$ and $\tilde{C}\geq 1/2$. The configuration $M=a\gamma$ where $\gamma^2=1$ with $\gamma$ containing an equal numbers of $+1$ and $-1$ is also a solution iff
\begin{eqnarray}
a^2=-\frac{\tilde{B}}{\sqrt{N}(2\tilde{C}+3)}({\rm non~uniform~ordered}).
\end{eqnarray}
By substituting the configurations $M=a{\bf 1}_{2N}$ and $M=a\gamma$ in the Schwinger-Dyson identity (with $V=V_0+V_1$) 
\begin{eqnarray}
\frac{<V>}{N^2}=\frac{1}{4}+\frac{B}{2}\frac{<Tr M^2>}{N^2}+\frac{B^{'}}{2} \frac{<(TrM)^2>}{N^2},
\end{eqnarray}
we obtain the energies $E_1$ and $E_2$ in the two phases. 
It is not difficult to show that $E_1\leq E_2$ iff (recall that $\tilde{B}+1\leq 0$, $2\tilde{C}-1\geq 0$)
\begin{eqnarray}
\tilde{B}+1\leq -\frac{2\tilde{C}-1}{4}.
\end{eqnarray}
This result is confirmed non-perturbatively using Monte Carlo simulation of the eigenvalue problem corresponding to the diagonalization of the quartic multitrace matrix model $V=V_0+V_1$. Thus, there exists a region in the phase diagram in which the uniform ordered configuration is more stable than the disordered and the non-uniform ordered configurations \cite{Ydri:2015vba,Ydri:2015zsa}.
 
The partition function of the theory is given by
\begin{eqnarray}
Z=\int {\cal D} M \exp(-V[M]).
\end{eqnarray}
We will assume now that the matrix $M$ is $2N\times 2N$. Then, without any loss of generality, we can expand the matrix $M$ as
\begin{eqnarray}
M=M_0{\bf 1}_{2N}+M_1~,~Tr M_1=0.
\end{eqnarray}
Hence
\begin{eqnarray}
M_1=\sigma_aX_a~,~
M_0=a+m,
\end{eqnarray}
where $\sigma_a$ are the standard Pauli matrices, $m$ is the fluctuation in the zero mode, and $X_a$ are three hermitian $N\times N$ matrices. By substitution, we obtain immediately the model
\begin{eqnarray}
Z&=&\int {\cal D} X_a \exp(-V[X]))\int dm\exp(-f[m]).\label{fundamental3}
\end{eqnarray}
The potential $V$ is given now by the $SO(3)-$symmetric three matrix model 
\begin{eqnarray}
V&=&-CTr[X_a,X_b]^2+2CTr(X_a^2)^2+4D(Tr X_a^2)^2\nonumber\\
&+&2(B+\beta_0 a^2)TrX_a^2+2ia\gamma\epsilon_{abc} TrX_aX_bX_c.\label{fundamental2}
\end{eqnarray} 
The new coefficients $\beta_0$ and $\gamma$ are given below. 

We observe that the Chern-Simons term is proportional to the value $a$ of the order parameter. Thus, it is non-zero only in the Ising phase, and as a consequence, by tuning the parameters appropriately to the region in the phase diagram where the Ising phase exists, we will induce a non-zero value for the Chern-Simons. This is effectively the Myers term responsible for the condensation of the geometry \cite{Myers:1999ps,Azuma:2004zq}. The above multitrace three matrix model is then precisely a random matrix theory describing non-commutative gauge theory on the fuzzy sphere, where the first term is the Yang-Mills piece, whereas the second and fourth terms combine to give mass and linear terms for the normal scalar field on the sphere (recall that $a$ runs from $1$ to $3$). The third doubletrace term, proportional to $D$, depends only on the zero mode of the normal scalar field. Thus, it is not expected to play a major role in our discussion here. If we simply set $D=0$ then we get essentially the random matrix theory describing non-commutative gauge theory on the fuzzy sphere  found in   \cite{Steinacker:2003sd} (notice that the parameter $C$ appears in front of the Yang-Mills as well as in front of the normal mass term and thus can be scaled away). This theory itself is a generalization of the stringy non-commutative gauge theory on the fuzzy sphere considered in \cite{Alekseev:2000fd,Iso:2001mg,CarowWatamura:1998jn}.

Furthermore, we have shown recently in \cite{Ydri:2016osu} that the above model with $D=0$ sustains an absolutely stable emergent fuzzy two-sphere geometry, and as a consequence, the expansion around this emergent geometry to obtain a non-commutative gauge theory is fully consistent for all values of the gauge coupling constant. In other words, there is no phase transition to a Yang-Mills matrix phase at a finite value of the gauge coupling constant. This is expected to hold also for $D\neq 0$. 

However, we should emphasis here that we have obtained dynamically this slightly generalized version of non-commutative gauge theory on the fuzzy sphere by going to the phase of the model where a non-zero uniform order persists, and then by expanding around this order, we secure a non-zero Chern-Simons term crucial for the underlying emergent geometry of the fuzzy sphere. Recall that in \cite{Steinacker:2003sd}, this was achieved by constraining the matrix $M$ directly in a particular way. 
 
What is the effect of the zero mode $m$?

The potential $f$ in (\ref{fundamental3}) of the zero mode $m$ is given by
\begin{eqnarray}
f
&=&\big[2i\gamma\epsilon_{abc} TrX_aX_bX_c+4\alpha a^3+4\beta_0 aTrX_a^2+2a\beta_1\big]m\nonumber\\
&+&\big[2\beta_0Tr X_a^2+\beta_1+6\alpha a^2\big] m^2+4\alpha a m^3+\alpha m^4.
\end{eqnarray}
The various new coefficients, for our case (\ref{our_case}), are given by

\begin{eqnarray}
&&\beta_0=3N^2(2\tilde{C}-1)~,~\beta_1=2N^{5/2}(\tilde{B}+1)\nonumber\\
&&\alpha=N^3(2\tilde{C}-1)~,~\gamma=2N^2(2\tilde{C}-1).
\end{eqnarray}
The integration over $m$ can be done. The leading contribution in the large $N$ limit is essentially the one-loop result and it is by construction subleading compared to (\ref{fundamental2}). This integral consists of some function of $TrX_a^2$ and $i\epsilon_{abc}TrX_aX_bX_c$. The conclusion of the foregoing discussion remains practically unchanged with the addition of these multitrace corrections.   


Thus, the multitrace one matrix model (\ref{fundamental1}), which involves only a single hermitian matrix $M$ with $U(2N)$ symmetry, and which exhibits a uniform order for some values of the mass parameter $B$ and the quartic coupling constant $C$,  provides a new mechanism for emergent fuzzy two-sphere and its non-commutative gauge theory. Indeed, by expanding around the uniform order, and then performing the integral over the zero mode $m$, it is seen that the resulting theory given by the mostly (since $D\neq0$) single trace $SO(3)-$symmetric three matrix model (\ref{fundamental2}), plus  the multitrace corrections induced by the integral over $m$, is a matrix model describing a non-commutative gauge theory on the fuzzy sphere. The multitrace corrections are all subleading in $N$, with the exception of the term $(TrX_a^2)^2$ in the model  (\ref{fundamental2}),  and they depend generically on the mass deformations $TrX_a^2$ and $i\epsilon_{abc}TrX_aX_bX_c$. 

The model (\ref{fundamental2}) with $D=0$ has been shown in \cite{Ydri:2016osu} to sustain an absolutely stable emergent fuzzy sphere with fluctuation given by a non-commutative $U(1)$ gauge theory very weakly coupled to a scalar field (playing the role of dark energy). In the full multitrace model there exists always the possibility of a phase transition  from the fuzzy two-sphere phase to a Yang-Mills matrix phase \footnote{In Yang-Mills matrix models with a non-zero Chern-Simons term there exists typically a phase transition from the fuzzy sphere phase to a Yang-Mills matrix phase.} which occurs at some finite value of the gauge coupling constant. However, this is not expected since $(TrX_a^2)^2$ depends only on the zero mode of the normal scalar field while all other multitrace corrections are subleading in $N$. In other words, the full multitrace matrix model is also expected to sustain an absolutely stable emergent fuzzy sphere with fluctuation given by a slightly generalized non-commutative $U(1)$ gauge theory coupled to a scalar field.

In summary, the multitrace matrix model  (\ref{fundamental1}) gives a novel and a much simpler  alternative to Connes noncommutative geometry \cite{Connes:1996gi} and to the IKKT matrix model \cite{Ishibashi:1996xs} for emergent geometry in two dimensions.

\paragraph{Acknowledgments:}
This research was supported by CNEPRU: "The National (Algerian) Commission for the Evaluation of
University Research Projects"  under contract number ${\rm DO} 11 20 13 00 09$. We would like also to acknowledge generous funding and warm hospitality of DIAS (Dublin) and ICTP (Trieste).

\end{document}